Give an equation whose solution set is the line joining (2,3) and(-4,6) What is its equation in slope-intercept form?
Answer by stanbon(75887)

Give an equation whose solution set is the line joining (2,3) and(-4,6) What is its equation in slope-intercept form?
----------------
slope = (6-3)/(-4-2) = 3/-6 = -1/2
-------------------------
Form: y = mx+b
3 = (-1/2)2 + b
---
3 = -1 + b
b = 4
-----
y = (-1/2)x+4
====================
Cheers,
Stan H.
====================